OpenStudy (anonymous):

use addition or sub to simplify sin θ cos 3θ + cos θ sin 3θ = 0

OpenStudy (anonymous):

Remember that \(\sin(x+y)=\sin(x)\cos(y)+\sin(y)\cos(x)\).

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

At least it looks like they want you to use that property.

OpenStudy (anonymous):

So identify what corresponds to what. What does x correspond to and what does y correspond to.

OpenStudy (anonymous):

xis theta y 3 theta

OpenStudy (anonymous):

So then what is \(\sin(x+y)\)?

OpenStudy (anonymous):

sin 4 theata

OpenStudy (anonymous):

???

OpenStudy (anonymous):

ya its \[\sin (4\theta)=0\]

OpenStudy (anonymous):

and on the interval of o to 2pi???

OpenStudy (anonymous):

what the values of theta on this interval

OpenStudy (anonymous):

i dont knoe

OpenStudy (anonymous):

i mean u need to find what in this interval

OpenStudy (anonymous):

i dont know how tho